If α and β are acute and tan α =
1
7
, and cot β =
1
7
, find:
(i) sin β

Answer by stanbon(75887)

If α and β are acute and tan α = 1/7
-------------------------
and cot β = 1/7
------
, find:
(i) sin β
---
By definition, cos = x/r
Since cos(B) = 1/7, x = 1 and r = 7
-----
Then y = sqrt[7^2-1^2] = 4sqrt(3)
-----
Ans: sin(B) = y/4 = 4sqrt(3)/7
-----------------------------------
Cheers,
Stan H.
------------------